A sequence is bounded if and only if it is both bounded above and bounded below


1
Expert's answer
2021-02-01T07:37:49-0500

Let "\{a_n\}_{n=1}^{\infty}" be a bounded sequence, that is there exists a constant M>0, such that "|a_n|\leq M" .

Then "a_n\leq M" and, hence, this sequence is bounded above.

We have also "a_n\geq -M" and, hence, this sequence is bounded below.


Let "\{a_n\}_{n=1}^{\infty}" be a sequence bounded above and bounded below.

Tthat is there exist two constants M1 and M2 such that

"M_1\leq a_n\leq M_2"

Let "M=\max\{|M_1|, |M_2|\}" . Then we have

"a_n\leq M_2\leq |M_2|\leq \max\{|M_1|, |M_2|\}= M" and

"a_n\geq M_1\geq -|M_1|\geq -\max\{|M_1|, |M_2|\}= -M"

Hence, "|a_n|\leq M" and the sequence "\{a_n\}_{n=1}^{\infty}" is bounded.